OpenStudy (anonymous):

In 2 years, Mark will be 1/2 the age of Alan. Alan is currently 28 years old. Which equation could be used to find Mark's current age (x)? A. 1/2 (x + 2) = (28 + 2) B. 1/2x + 2 = (28 + 2) C. (x + 2) = 1/2 (28 + 2) D. (x + 2) = 1/2 (28) + 2

OpenStudy (cwrw238):

C

OpenStudy (campbell_st):

Mark's age is x in 2 years x+ 2 Allan's age 28 in to years 28 + 2 Mark will be 1/2 allan's x + 2 = 1/2(28+2)

OpenStudy (anonymous):

x+2=1/2(28+2) x=13